A flat square piece of cardboard is used to construct an open box. Cutting a 3 feet by 3 feet square off of each corner and folding up the edges will yield an open box (assuming these edges are taped together). If the desired volume of the box is 147 cubic feet, what are the dimensions of the original square piece of cardboard?
Answer by mananth(16946)

Let the square card board have length = x
#feet is cut from each corrner
so the length decreases by 6 feet.
new length = (x-6)
height = 3 ft.
Volume = (x-6)*(x-6)*3
(x^2-12x+36)*3=147
/3
x^2-12x+36 = 49
x^2-12x-13=0
x^2-13x+x-13=0
x(x-13)+1(x-13)=0
(x-13)(x+!)=0
x=13 OR -1
ignore negative
the length is 13 feet